Find all solutions of the equation in the interval [0,2π)
3 cot^2 x-9 = 0

Found 2 solutions by Alan3354, Zakia B.:
Answer by Alan3354(69443)

Find all solutions of the equation in the interval [0,2π)
3 cot^2 x-9 = 0
cot^2(x) = 3
tan^2(x) = 1/3
tan(x) = +sqrt(3)/3
x = pi/6, 7pi/6
-----
tan(x) = -sqrt(3)/3
x = 5pi/6, 11pi/6

Answer by Zakia B.(1)

3 cot^2 x-9 = 0
cot^2(x) = 3
tan^2(x) = 1/3
tan(x) = +sqrt(3)/3
x = pi/6, 7pi/6
-----
tan(x) = -sqrt(3)/3
x = 5pi/6, 11pi/6